If we let only two forces acting on two  bodies be their mutual interactions . And both bodies start from rest, how do we show the distance traveled by each is inversely proportional to its mass ?

Grade:12th pass

1 Answers

ROSHAN MUJEEB
3 years ago
Dear Student
When two bodies are coming closer only due to gravitational force, the momentum is conserved. Let M1, m2 be the bodies and v1 and V2 the respective velocities.
M1V1= m2v2
M1/m2 = v2/V1
also V2/V1= D2 /D1 , Distance is proportional to velocity
Therefire M1/m2= D2/D1
hope you got it